# Extending Audio Context for Long-Form Understanding in Large Audio-Language Models

### v1

@@ 6.1 Methodology
The goal of VLAT is to simulate and expose models to audio context windows of diverse lengths during training. Let $L_{\text{data}}$ be the actual length of a training audio sample. For each sample, we obtain a “virtual” source length, $L_{\text{virt}}$, as the model’s default audio context length times a factor randomly sampled from the range [1, 5, 10, 15, 20, 25]x. We then apply Partial YaRN [^6] Whole YaRN is not used here, as we observe that it diverges quickly when training under the VLAT framework. to stretch or compress the $L_{\text{virt}}$-long positional window to $L_{\text{data}}$.
@@
Randomizing positional encoding has previously been explored in Ruoss et al. (2023), where positional indices are randomly downsampled to let the model see larger positional values. Rather than using $\{1,2,\dots,N\}$ for an length-$N$ input, they use a randomly subsampled values such as $\{1,4,\dots,L\}$ with $L>N$ instead. In contrast, VLAT differs in employing interpolation inside the context window to create a dense, continuous space of positions, unlike sparse integer subsampling. It is also bidirectional, teaching the model through both compressed and stretched contexts. Furthermore, VLAT is a targeted, modality-bound method that modifies only the audio tokens, preserving the base LLM’s original textual space.
@@ 6.2 Evaluation
We fine-tune Qwen2-Audio on a 2mins YODAS2-MCQA training set [^7] To prevent data leakage, the train/test splits were performed at the video level, ensuring no source material overlaps between the training set and any of the test sets. using two methods: a standard Vanilla fine-tuning and VLAT. We employ 10-epoch qkvo LoRA for the adaptations. We then evaluate the models from both training methods under two different extensions at inference: Vanilla and Partial PI. See Appendix E.1 for additional detail on VLAT configuration.
@@ 6.3 Results
Furthermore, as can be seen in "Partial PI Inference" columns, VLAT’s benefit is complementary to inference-time extension. When Partial PI is applied during evaluation, the VLAT-trained model again outperforms its vanilla-trained counterpart, achieving the highest overall 10mins performance of 81.73%. This indicates that VLAT and inference-time extension are compatible strategies, and their combination yields the most robust long-context performance.
@@ 7.2 Partial YaRN with Three-group Frequency Partitioning
For the Partial YaRN method, we modify the original 3-group frequency partition from original YaRN to a 2-group method as discussed in Section 3.1 and Appendix B. Here we conduct a direct empirical comparison between the 2-group and 3-group approaches.
@@
We see that our 2-group partitioning broadly outperform 3-group in our modality-specific scenario. This is most evident with the Qwen2-Audio model, where the 3-group method suffers a large performance drop on 5mins and 10mins audio, with an accuracy gap of 28.53% and 12.00% respectively. Overall, this result validates the effectiveness and reliability of our proposed 2-group scheme used in Partial YaRN.
8 Conclusion
@@
This work addressed the challenge of extending audio context windows of LALMs by studying the application of LLM’s whole-context extension methods, and proposing a training-free audio-only extension Partial YaRN and the VLAT training strategy. We demonstrate the effectiveness of extension-based methods over non-extension baselines across various long-audio settings. Leveraging our finding that SALMONN and Qwen2-Audio retain consistent performance up to 2mins audio window, we achieved an even stronger performance by extending from the 2mins window instead of 30s. Later, we showed that fine-tuning through VLAT helps the models to generalize to audio contexts far exceeding the length of the training data, offering a robust and data-efficient pathway to develop LALMs with better longform understanding. Future work could explore a similar modality-bound extension strategy for video models.
@@
Our experiments are conducted exclusively on the multiple-choice question answering (MCQA) tasks. We made this choice to ensure a precise and unambiguous measurement of the model performance under different audio lengths. This mitigates the confounding variables inherent in evaluating open-ended generation, such as judging the semantic equivalence or stylistic alignment. However, a key tradeoff of this focused evaluation is its limited ability to assess the nuanced generative and language modeling capabilities of the base LLM. Consequently, while our results validate the model’s ability to retrieve information from audio, they do not fully test the importance of text preservation of the audio-only methods like Partial YaRN. Furthermore, our study does not assess performance across other long-audio task formats found in benchmarks such as (Guo et al., 2025; Ahia et al., 2025; Goel et al., 2025).
@@
Unlike the original YaRN which has generally good predefined cutoff dimensions and a closed-form formula for determining the attention temperature, our proposed Partial YaRN requires tuning of these two hyperparameters, thereby introducing additional computational burden.
@@ Acknowledgments
YC would like to thank Ekapol Chuangsuwanich for insightful discussion on the earlier version of this work. He also thanks Thanapat Trachu for valuable guidance on the illustration of Partial YaRN.
Appendix A Extended Related Works
Positional encodings.
Conceived along with the introduction of transformers (Vaswani et al., 2017), the original absolute positional encoding utilizes sinusoidal waves of varying frequencies to uniquely represent absolute position of the tokens, and is incorporated into the model by adding directly into the input sequence. Later works (Devlin et al., 2019; Lan et al., 2020; Clark et al., 2020) propose to use trainable vectors for representing absolute positional information, where token at each position is summed with a learned vector, allowing the models to learn the optimal positional encoding for themselves. However, it cannot extrapolate to sequences longer than the maximum length it was trained on. Shaw et al. (2018) introduces to encode the relative position of the token instead, where the relative distance between the query and key is captured and injected into every attention layer. Dai et al. (2019) modifies the decomposed form of the attention equation, and introduces additional trainable location-agnostic key vectors to distinguish between content-based and location-based querying behaviors. Raffel et al. (2020) groups the relative distances into buckets, and associate each bucket with a learnable scalar for adding into the attention logits. Simplifying the encoding scheme and reducing trainable parameters. TUPE (Ke et al., 2021) uses separate linear projections for the word and positional information. ALiBi (Press et al., 2022) simply adds linear biases with a learnable slope into the attention scores, the bias grows linearly with the relative distance between the tokens. This method generally imposes a strong locality preference by encouraging models to attend more to the nearby tokens, while discouraging them from attending to tokens that are farther away. Several later works extend RoPE for multimodal modeling (e.g. Multimodal RoPE Bai et al. (2025), VideoRoPE (Wei et al., 2025), VRoPE (Liu et al., 2025)).
Large video-language models.
Video input is similar to the audio modality in that they’re both naturally continuous and can be of arbitrary length. Many Large Video-Language Models (LVLMs), such as Video-LLaMA (Zhang et al., 2023b) and Video-ChatGPT (Maaz et al., 2024), adopt architectural paradigm similar to the unified input space LALMs we study, where video frames are embedded and projected into a base LLM’s input space. However, the context length problem is even more exacerbated in the video domain due to the high dimensionality and large number of tokens needed to represent even short videos. To improve efficiency, the LVLM community has largely focused on input compression techniques like query aggregation (Zhang et al., 2023b; Li et al., 2023a; Song et al., 2024), frame pooling (Luo et al., 2023; Maaz et al., 2024), and feature merging (Weng et al., 2024). These strategies primarily aim to reduce the number of tokens before they enter the LLM. Our work offers a complementary perspective: instead of altering the input, we manipulate the positional space within the model to accommodate a longer sequence. The principles of Partial YaRN and Virtual Longform Audio Training could therefore be potentially applicable to the video domain as well, hopefully providing an orthogonal path to achieving long-context understanding in synergy with existing compression methods.
Appendix B Elaboration on Two-group Partitioning of Partial YaRN
@@
See Appendix 7.2 for direct performance comparison between 2 and 3-group frequency partitioning versions of Partial YaRN.
Appendix C Implementation of Partial YaRN via position_ids Manipulation
A key implementation detail of our work is how we apply the positional stretching to only a subsection of the input sequence. While the original YaRN implementation modifies the RoPE frequency variables ($\boldsymbol{\theta}$), this approach isn’t straightforward for a partial modification. Instead, our implementation of Partial YaRN directly manipulates the position_ids and multiply it with the unmodified frequencies. This section details the implementation our method.
@@ D.1 Training and Inference
We use LoRA (Hu et al., 2022) of rank 8 for fine-tuning through HuggingFace’s PEFT library (Mangrulkar et al., 2022). With alpha set to 16, and a small LoRA dropout of 0.5%. Models are trained with AdamW optimizer (Loshchilov and Hutter, 2019), learning rate is set to 5e-5, a batch size of 8 samples, and gradient norm clipped to 1.0 (Pascanu et al., 2013).
@@ D.3 Dataset Construction
We specifically use a portion the en129888https://huggingface.co/datasets/espnet/yodas2/tree/main/data/en129 subset of the YODAS2 dataset (Li et al., 2023b) (CC-BY-3.0) through HuggingFace’s Datasets library (Lhoest et al., 2021). The dataset is then split into train, validation, and test sets. We guarantee no test leakage in a way that if a dialogue is in the test set, then all the other dialogues from the same video are enforced to be in the test set as well.
YODAS2 MCQA inference prompt Your answer MUST be in the format of "(NUMBER) STATEMENT". For example, if the answer was (4) A pen, you would ONLY output "(4) A pen". Do NOT include any other text. Using information from all the given audios. {question}
Figure 2: Prompt used for inferencing on YODAS2-MCQA.
To obtain specific audio length, we chunk the long audio in the dataset to 1, 2, 5, and 10 minutes. Last, incomplete chunk of each long-audio is also included as well. Each of the chunk will be used to generate 5 MCQA pairs via Gemini 2.0 Flash (Team et al., 2023). We provide the full MCQA generation prompt in Figure 7.
@@ E.1 Different VLAT’s Virtual Length Sampling Strategies
Analysis.
Given this uncertainty, we retained the Default strategy for our main experiments in Section 6. Its factor range intuitively and completely covers the lengths seen during evaluation, making it a more principled and safer choice.
@@ E.2.2 Sensitivity on SALMONN
YODAS2 MCQA generation prompt You are an expert in creating high-quality Multiple-Choice Question Answering (MCQA) datasets for speech understanding. Your task is to generate QA pairs based on audio segments. **Input:** You will be given a segment of an audio recording. **Task:** For each given transcript, generate **5 distinct 4-Choice Question-Answer (QA) pairs** that each has 1 correct and 3 incorrect choices (so four choices in total), and are factually grounded in the content of the transcript. The questions should cover different aspects of the audio segment, such as: * **Who:** Identify speakers or people mentioned. * **What:** Describe actions, objects, or events. * **When:** Determine times or dates mentioned. * **Where:** Locate places or locations. * **Why/How:** Explain reasons or processes. * **Details:** Ask for specific details mentioned in the audio segment. * **Purpose/Intention:** Infer the speaker’s intent. * **Summary:** Condense information from the audio segment. * **Count:** How many people, men or women in the audio segment. * **Opinion/Sentiment:** What opinions or sentiment is expressed. * **Negation:** Understand what is NOT being said or done. **Output:** Your output should be a JSON formatted string. The JSON should have the following structure: "response": [ "question": "Question 1", "correct": "Correct answer", "incorrect1": "Incorrect answer1", "incorrect2": "Incorrect answer2", "incorrect3": "Incorrect answer3", "question": "Question 2", "correct": "Correct answer", "incorrect1": "Incorrect answer1", "incorrect2": "Incorrect answer2", "incorrect3": "Incorrect answer3", "question": "Question 3", "correct": "Correct answer", "incorrect1": "Incorrect answer1", "incorrect2": "Incorrect answer2", "incorrect3": "Incorrect answer3", "question": "Question 4", "correct": "Correct answer", "incorrect1": "Incorrect answer1", "incorrect2": "Incorrect answer2", "incorrect3": "Incorrect answer3", "question": "Question 5", "correct": "Correct answer", "incorrect1": "Incorrect answer1", "incorrect2": "Incorrect answer2", "incorrect3": "Incorrect answer3" ] Important Guidelines: Accuracy: Ensure the answers are directly supported by content of the audio. Do not introduce information that is not present in the audio. Relevance: The questions should be relevant to the content of the audio segment. Diversity: The 5 QA pairs should be diverse in the type of question asked (who, what, when, where, why, how, detail, etc.). Try to use each question type at least once, if applicable to the transcript. Clarity: The questions and answers should be clear, concise, and grammatically correct. Completeness: The answers should fully address the question and provide sufficient information. Conciseness: The answers should be as concise as possible while still being accurate and complete. Coverage: The questions should collectively cover the entire audio segment. Aim for diverse coverage for each question. Prioritize the most salient and important information. Avoid Ambiguity: Ensure the questions are not ambiguous and have a clear, single answer. Plausibility of Incorrect Answers: The incorrect answers should be plausible but wrong, and related to the context of the transcript. Don’t make them obviously incorrect or nonsensical. Use information or wording from the transcript to make them believable distractors. JSON Validity: The output must be a valid JSON string. Double-check the formatting and escaping of special characters. Context-Awareness: While each audio segment should be treated independently, the types of questions should logically relate to the context of human conversations. Try to create questions which someone would naturally ask. Answer Length Consistency: Try to make the length of the correct and incorrect answers as similar as possible. This makes it more difficult to guess the answer based on length alone. Example: Input Transcript: "Okay, so the meeting is scheduled for next Tuesday at 2 PM in Conference Room A. We’$l$ be discussing the Q3 marketing report and the proposed budget changes." Output: "response": [ "question": "When is the meeting scheduled?", "correct": "Next Tuesday at 2 PM", "incorrect1": "This Friday at 10 AM", "incorrect2": "Next Wednesday at 3 PM", "incorrect3": "Today at 4 PM" , "question": "Where will the meeting take place?", "correct": "Conference Room A", "incorrect1": "Conference Room B", "incorrect2": "The Main Auditorium", "incorrect3": "The Executive Boardroom" , "question": "What marketing report will be discussed?", "correct": "The Q3 marketing report", "incorrect1": "The Q2 marketing report", "incorrect2": "The annual marketing report", "incorrect3": "The preliminary marketing report" , "question": "Besides the marketing report, what else will be discussed?", "correct": "The proposed budget changes", "incorrect1": "The sales figures", "incorrect2": "The employee satisfaction survey", "incorrect3": "The new product launch" , "question": "What is the purpose of the meeting?", "correct": "To discuss the Q3 marketing report and budget changes", "incorrect1": "To plan the annual company picnic", "incorrect2": "To interview new job candidates", "incorrect3": "To celebrate the company’s anniversary" ] Now, generate 5 QA pairs in the specified JSON format for the input audio:

### v2

@@ 6.1 Methodology
The goal of VLAT is to simulate and expose models to audio context windows of diverse lengths during training. Let $L_{\text{data}}$ be the actual length of a training audio sample. For each sample, we obtain a “virtual” source length, $L_{\text{virt}}$, as the model’s default audio context length times a factor randomly sampled from the range [1, 5, 10, 15, 20, 25]x. We then apply Partial YaRN [^7] Whole YaRN can’t be used here, as we observe that it diverges quickly when training under the VLAT framework. to stretch or compress the $L_{\text{virt}}$-long positional window to $L_{\text{data}}$.
@@
Randomizing positional encoding has previously been explored in Ruoss et al. (2023), where positional indices are randomly downsampled to let the model see larger positional values. Rather than using $\{1,2,\dots,N\}$ for an length-$N$ input, they use a randomly subsampled values such as $\{1,4,\dots,L\}$ with $L>N$ instead. In contrast, VLAT differs in employing interpolation inside the context window to create a dense, continuous space of positions, unlike sparse integer subsampling. It is also bidirectional, teaching the model through both compressed and stretched contexts. Furthermore, VLAT is a targeted, modality-decoupled method that modifies only the audio tokens, preserving the base LLM’s original textual space.
@@ 6.2 Evaluation
We fine-tune Qwen2-Audio on the 2mins YODAS2-MCQA training set [^8] To prevent data leakage, the train/test splits were performed at the video level, ensuring no source material overlaps between the training set and any of the test sets. using two methods: a standard Vanilla fine-tuning, and VLAT. We employ 10-epoch qkvo LoRA. We then evaluate the models from both training methods under two different extensions at inference: Vanilla and Partial PI. Results are averaged from 3 random seeds. See Appendix E.1 for detail on VLAT configuration.
@@ 6.3 Results
Furthermore, as can be seen in "Partial PI Inference" columns, VLAT’s benefit is complementary to inference-time extension. When Partial PI is applied during evaluation, the VLAT-trained model again outperforms its vanilla-trained counterpart, achieving the highest overall 10mins performance of 81.73%. This indicates that VLAT and inference-time extension are compatible strategies, and their combination yields the most robust long-audio-context performance.
We show in Appendix E.1 that VLAT’s performance can be even more enhanced by using a narrower range for virtual length $L_{\text{virt}}$.
@@ 7.2 Partial YaRN with Three-group Frequency Partitioning
For the Partial YaRN method, we modify the original 3-group frequency partition from original YaRN to a 2-group method as discussed in Section 3.1 and Appendix A. Here we conduct a direct empirical comparison between the 2-group and 3-group approaches.
@@
We see that our 2-group partitioning broadly outperform 3-group in our modality-specific scenario. This is most evident with the Qwen2-Audio model, where the 3-group method suffers a large performance drop on 5mins and 10mins audio, with an accuracy gap of 28.53% and 12.00% respectively. Overall, this result validates the effectiveness and reliability of our proposed 2-group scheme.
8 Related Work
Positional encodings.
Absolute positional encoding was originally proposed in Vaswani et al. (2017). It utilizes sinusoidal waves to uniquely represent absolute position of each token. Later, (Devlin et al., 2019; Lan et al., 2020; Clark et al., 2020) use trainable vectors for representing absolute positions, allowing the models to learn the positional encoding for themselves. However, it cannot extrapolate to sequences longer than the maximum length it was trained on. Shaw et al. (2018) encodes the relative distances between the tokens instead of absolute positions. Dai et al. (2019) introduces additional trainable position-agnostic components to distinguish between content-based and location-based attention behaviors. Raffel et al. (2020) groups the relative distances into buckets, and associate each bucket with a learnable scalar for adding to attention logits. TUPE (Ke et al., 2021) uses separate linear projections for the token and positional encodings. ALiBi (Press et al., 2022) simply adds linear biases with a learnable slope into the attention scores, imposing a strong locality preference. Several later works extend RoPE for multimodal modeling, such as Multimodal RoPE Bai et al. (2025), VideoRoPE (Wei et al., 2025), and VRoPE (Liu et al., 2025b).
Large video-language models.
Video input is similar to the audio in that they’re both naturally continuous and can be of arbitrary lengths. Many Large Video-Language Models (LVLMs) (Zhang et al., 2023b; Maaz et al., 2024; Lin et al., 2024; Jin et al., 2024; Ren et al., 2024), adopt architectural paradigm similar to the unified input space LALMs we studied, where video frames are embedded and projected into a base LLM’s input space. However, the context length problem is even more exacerbated in the video domain due to the high dimensionality and large number of tokens needed to represent even short videos. To improve efficiency, the LVLM community has largely focused on input compression techniques like query aggregation (Zhang et al., 2023b; Li et al., 2025a; Song et al., 2024), frame pooling (Luo et al., 2023; Maaz et al., 2024), and feature merging (Weng et al., 2024). These strategies primarily aim to reduce the number of tokens before they enter the base LLM. Our work offers a complementary perspective: instead of altering the input, we manipulate the positional space within the model to accommodate a longer sequence, similar to Zhang et al. (2025) in LVLMs. The principles of Partial YaRN and VLAT are likely applicable to the video domain as well, hopefully providing a distinct path to achieve long-video-context understanding and generative control in synergy with existing methods.
Long-audio LALMs.
Concurrent works have proposed LALMs that have inherently long audio-context. Audio Flamingo 3 (Ghosh et al., 2025a) is capable of processing audio up to 10mins due long-audio training. Voxtral (Liu et al., 2025a) can handle up to 40mins due to the long 32k context of its LLM backbone. FastLongSpeech (Guo et al., 2025) proposes Iterative Fusion which incrementally merges the audio representation based on their information density. This compressed representation facilitates efficient processing of long audio. Qwen3-Omni (Xu et al., 2025) is a multimodal model that can accept text, video, and audio as input. It employs a custom Audio Transformer which is trained in an encoder-decoder fashion similar to the Whisper model. Qwen3-Omni is also capable of processing audio of up to 40mins.
While these models achieve impressive long-audio performance, they do so by designing new architectures, curating large volumes of long-audio data, and undertaking extensive training. Our work investigates an orthogonal yet complementary direction of efficiently extending the context of existing LALMs. Rather than requiring costly data acquisition and retraining, we provide a lightweight, architecture-agnostic "drop-in" enhancement (Partial YaRN) and an efficient fine-tuning strategy (VLAT), offering a practical and accessible approaches for improving the long-audio capabilities of a wide range of existing models.
9 Conclusion and Future Work
@@
This work addressed the challenge of extending audio context windows of LALMs by studying the application of LLM’s whole-context extension methods, and proposing a training-free audio-only extension Partial YaRN and the VLAT training strategy. We demonstrated the effectiveness of extension-based methods over non-extension baselines across various long-audio settings. Leveraging our finding that SALMONN and Qwen2-Audio retain consistent performance up to 2mins audio window, we achieved an even stronger performance by extending from the 2mins window instead of 30s. Later, we showed that fine-tuning through VLAT allows the models to generalize to audio contexts far exceeding the lengths of the training data, offering a robust and data-efficient pathway to develop LALMs with better longform understanding.
Future work could apply Partial YaRN and VLAT to vision models. Beyond content understanding, these approaches could also be explored for controllable generation in multimodal models, such as multi-scale image synthesis, or even more finegrained decoupling for non-uniform speed control throughout audio or video generation.
@@
Our experiments are conducted exclusively on the multiple-choice question answering tasks. We made this choice to ensure a precise and unambiguous measurement of the model performance under different audio lengths. This mitigates the confounding variables inherent in evaluating open-ended generation, such as judging the semantic equivalence or stylistic alignment. However, a key tradeoff of this focused evaluation is its limited ability to assess the nuanced generative and language modeling capabilities of the base LLM. Consequently, while our results validate the model’s ability to retrieve information from audio, they do not fully test the importance of text preservation of the audio-only methods like Partial YaRN. Furthermore, our study does not assess the performance across other long-audio task formats found in benchmarks such as Guo et al. (2025), Ahia et al. (2025), or Ghosh et al. (2025a).
Reliance on synthetic data.
Our study utilizes YODAS2-MCQA, a custom dataset synthesized from the YODAS2 corpus using Gemini 2.0 Flash. While we carefully steered the synthesis process and validated a subset of the data, it’s still possible for the synthesized samples to carry noise, repetition, and hallucinations that may affect the fidelity of our performance estimates compared to real-world scenarios.
@@
Unlike the original YaRN which has generally good predefined cutoff dimensions and a closed-form formula for determining the attention temperature, our proposed Partial YaRN requires tuning of two hyperparameters, thereby introducing additional computational burden.
@@ Acknowledgments
YC would like to thank Ekapol Chuangsuwanich for insightful discussion on the earlier version of this work. He also thanks Thanapat Trachu for valuable guidance on the illustration of Partial YaRN, and Konpat Preechakul for illuminating discussion on extensions of this work to the vision domain. We gratefully acknowledge the compute resources provided by Hatari NEXT, on which the experiments for this research were run.
Appendix A Elaboration on Two-group Partitioning of Partial YaRN
@@
See Section 7.2 for direct performance comparison between 2 and 3-group frequency partitioning versions of Partial YaRN.
Appendix B Implementation of Partial YaRN via position_ids Manipulation
A key implementation detail of our work is how we apply the positional stretching to only a subsection of the input sequence. While the original YaRN implementation modifies RoPE’s frequencies ($\boldsymbol{\theta}$), this approach isn’t straightforward for a partial modification. Instead, our implementation of Partial YaRN directly manipulates the position_ids and multiply it with the unmodified frequencies. This section details the implementation our method.
@@ Appendix B Implementation of Partial YaRN via position_ids Manipulation
Appendix C On Deriving Closed-Form Hyperparameter Formulae
A practical extension of this work is to derive closed-form formulae for determining the optimal cutoff dimension index and attention temperature for Partial YaRN akin to the solution provided in YaRN for unimodal LLMs (Peng et al., 2024), which would eliminate the cost of hyperparameter searches. We leave this for future work due to a key challenge not present in unimodal LLMs: the wide variance in audio representation lengths across different LALMs.
Unlike the relatively consistent token-to-information ratio in text across modern tokenizers, LALMs compress audio into discrete tokens at vastly different rates. For instance, a 30-second audio clip gets compressed into 88 tokens in SALMONN but 750 tokens in Qwen2-Audio. As our sensitivity analysis in Appendix E.2 shows, this disparity strongly influences the optimal hyperparameter choice, making closed-form formula non-trivial to derive.
@@ D.1 Training and Inference
We use LoRA (Hu et al., 2022) of rank 8 for fine-tuning through HuggingFace’s PEFT library (Mangrulkar et al., 2022). With alpha set to 16, and a small LoRA dropout of 5%. Models are trained with AdamW optimizer (Loshchilov and Hutter, 2019), learning rate is set to 5e-5, a batch size of 8 samples, and gradient norm clipped to 1.0 (Pascanu et al., 2013).
@@ D.2 Partial YaRN’s Hyperparameter Tuning
Figure 2: Distributions of audio segment lengths in the YODAS2-MCQA dataset, shown with logarithmic counts for improved visibility of the long-tail portions.
Figure 3: Distributions of correct answer choices for all question-answer pairs in the YODAS2-MCQA dataset.
YODAS2-MCQA inference prompt Your answer MUST be in the format of "(NUMBER) STATEMENT". For example, if the answer was (4) A pen, you would ONLY output "(4) A pen". Do NOT include any other text. Using information from all the given audios. {question}
Figure 4: Prompt used for inferencing on YODAS2-MCQA.
@@ D.3 Dataset Construction
Our dataset is derived from a portion of the en129999https://huggingface.co/datasets/espnet/yodas2/tree/main/data/en129 subset of YODAS2 (Li et al., 2023), a large-scale, multilingual corpus of YouTube-derived audio and speech data. The YODAS2 dataset is made available under the CC-BY-3.0 license. We accessed and processed the data using HuggingFace’s Datasets library (Lhoest et al., 2021).
To prevent video source leakage in the VLAT experiment, train-test split was performed at the video level; all dialogue segments from a single video were exclusively assigned to either the training set (2mins), or the test sets.
To obtain specific audio length, we chunk the long audio in the dataset to 1, 2, 5, and 10 minutes. Any remaining, incomplete chunk of each long-audio is also retained. Subsequently, we employed Gemini 2.0 Flash (Team et al., 2023) to generate five unique multiple-choice question-answer (MCQA) pairs for each audio chunk. The complete prompt utilized for this generation process is provided in Figure 9. The resulting dataset, which we have named YODAS2-MCQA, is publicly available at: https://huggingface.co/datasets/yophis/yodas2-mcqa.
For a more granular analysis of the YODAS2-MCQA dataset, we provide visualizations of the distributions of key characteristics. The distributions of audio segment lengths are presented in Figure 2. Additionally, the distribution of correct answers across the multiple-choice options is presented in Figure 3.
@@ E.1 Different VLAT’s Virtual Length Sampling Strategies
Analysis.
Given this uncertainty, we retained the Default strategy for our main experiments in Section 6. Its factors ranges intuitively and completely covers the lengths seen during evaluation, making it a more principled and safer choice.
@@ E.2.2 Sensitivity on SALMONN
YODAS2-MCQA generation prompt You are an expert in creating high-quality Multiple-Choice Question Answering (MCQA) datasets for speech understanding. Your task is to generate QA pairs based on audio segments. **Input:** You will be given a segment of an audio recording. **Task:** For each given transcript, generate **5 distinct 4-Choice Question-Answer (QA) pairs** that each has 1 correct and 3 incorrect choices (so four choices in total), and are factually grounded in the content of the transcript. The questions should cover different aspects of the audio segment, such as: * **Who:** Identify speakers or people mentioned. * **What:** Describe actions, objects, or events. * **When:** Determine times or dates mentioned. * **Where:** Locate places or locations. * **Why/How:** Explain reasons or processes. * **Details:** Ask for specific details mentioned in the audio segment. * **Purpose/Intention:** Infer the speaker’s intent. * **Summary:** Condense information from the audio segment. * **Count:** How many people, men or women in the audio segment. * **Opinion/Sentiment:** What opinions or sentiment is expressed. * **Negation:** Understand what is NOT being said or done. **Output:** Your output should be a JSON formatted string. The JSON should have the following structure: "response": [ "question": "Question 1", "correct": "Correct answer", "incorrect1": "Incorrect answer1", "incorrect2": "Incorrect answer2", "incorrect3": "Incorrect answer3", "question": "Question 2", "correct": "Correct answer", "incorrect1": "Incorrect answer1", "incorrect2": "Incorrect answer2", "incorrect3": "Incorrect answer3", "question": "Question 3", "correct": "Correct answer", "incorrect1": "Incorrect answer1", "incorrect2": "Incorrect answer2", "incorrect3": "Incorrect answer3", "question": "Question 4", "correct": "Correct answer", "incorrect1": "Incorrect answer1", "incorrect2": "Incorrect answer2", "incorrect3": "Incorrect answer3", "question": "Question 5", "correct": "Correct answer", "incorrect1": "Incorrect answer1", "incorrect2": "Incorrect answer2", "incorrect3": "Incorrect answer3" ] Important Guidelines: Accuracy: Ensure the answers are directly supported by content of the audio. Do not introduce information that is not present in the audio. Relevance: The questions should be relevant to the content of the audio segment. Diversity: The 5 QA pairs should be diverse in the type of question asked (who, what, when, where, why, how, detail, etc.). Try to use each question type at least once, if applicable to the transcript. Clarity: The questions and answers should be clear, concise, and grammatically correct. Completeness: The answers should fully address the question and provide sufficient information. Conciseness: The answers should be as concise as possible while still being accurate and complete. Coverage: The questions should collectively cover the entire audio segment. Aim for diverse coverage for each question. Prioritize the most salient and important information. Avoid Ambiguity: Ensure the questions are not ambiguous and have a clear, single answer. Plausibility of Incorrect Answers: The incorrect answers should be plausible but wrong, and related to the context of the transcript. Don’t make them obviously incorrect or nonsensical. Use information or wording from the transcript to make them believable distractors. JSON Validity: The output must be a valid JSON string. Double-check the formatting and escaping of special characters. Context-Awareness: While each audio segment should be treated independently, the types of questions should logically relate to the context of human conversations. Try to create questions which someone would naturally ask. Answer Length Consistency: Try to make the length of the correct and incorrect answers as similar as possible. This makes it more difficult to guess the answer based on length alone. Example: Input Transcript: "Okay, so the meeting is scheduled for next Tuesday at 2 PM in Conference Room A. We’$l$ be discussing the Q3 marketing report and the proposed budget changes." Output: "response": [ "question": "When is the meeting scheduled?", "correct": "Next Tuesday at 2 PM", "incorrect1": "This Friday at 10 AM", "incorrect2": "Next Wednesday at 3 PM", "incorrect3": "Today at 4 PM" , "question": "Where will the meeting take place?", "correct": "Conference Room A", "incorrect1": "Conference Room B", "incorrect2": "The Main Auditorium", "incorrect3": "The Executive Boardroom" , "question": "What marketing report will be discussed?", "correct": "The Q3 marketing report", "incorrect1": "The Q2 marketing report", "incorrect2": "The annual marketing report", "incorrect3": "The preliminary marketing report" , "question": "Besides the marketing report, what else will be discussed?", "correct": "The proposed budget changes", "incorrect1": "The sales figures", "incorrect2": "The employee satisfaction survey", "incorrect3": "The new product launch" , "question": "What is the purpose of the meeting?", "correct": "To discuss the Q3 marketing report and budget changes", "incorrect1": "To plan the annual company picnic", "incorrect2": "To interview new job candidates", "incorrect3": "To celebrate the company’s anniversary" ] Now, generate 5 QA pairs in the specified JSON format for the input audio: